the precision scientific instrument company manufactures thermometers that are supposed to give readings of 0 degrees C at the freezing point of water. Tests on alarge sample of these thermometers reveal that at the freezing point of water, some give readings below 0 degrees C(denoted by negative numbers) and some give readingsabove 0 degrees C (denoted by positive numbers). assume that the mean reading is 0 degrees C and the standard deviation of the readings is 1.00 degrees C. Also assumethat the frequency distribution errors closely resembles the normal distribution. A thermometer is randomly selected and tested. Find the temperature readingcorresponding to the given information.
Find the P96, the 96th percentile
Answer #1

μ=0
σ=1
Q3 means third quartile ,So P=75% =0.75

z score for P=0.75 is 0.67

x=z*σ + μ
=0.67*1 + 0
=0.67+0
=0.67